OpenStudy (anonymous):

A plane travels 2187 miles in 5.4 hours. Find the plane's speed in miles per hour. mi/h ?

OpenStudy (texaschic101):

2187/5.4 = x/1 -- (2187 miles to 5.4 hrs = x miles to 1 hr) cross multiply (5.4)(x) = (2187)(1) 5.4x = 2187 -- divide both sides by 5.4 x = 2187/5.4 x = 405 planes speed = 405 miles/hr

OpenStudy (anonymous):

thanks so much :)

OpenStudy (texaschic101):

anytime :)